OpenStudy (anonymous):

help?? Composite functions: If F(x)=3x+1 and g(x)=xˆ2+2x, find x when F(g(x)) = 10

OpenStudy (amoodarya):

F(x)=3x+1 g(x)=xˆ2+2x f(g(x))=f(x^2+2x) =3(x^2+2x)+1 now 3(x^2+2x)+1=10 solve for x 3(x^2+2x)=9 simplify by 3 and solve for x

OpenStudy (amoodarya):

(x^2+2x)=3 x^2+2x-3=0 (x-1) (x+3) =0 -1+3=2 -1*3=-3 can you go on ?

OpenStudy (anonymous):

um I'm confused at the part where you put -1+3=2

OpenStudy (amoodarya):

(x-a)(x-b)= x^2 -(a+b)x +(-a)(-b)

OpenStudy (anonymous):

oooh yeah ok so my final answer is -3?